Monika covers a distance of 10 km in 10 min.Rumpa covers a distance of 20 km in 12 min. Who runs faster??

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
speed of monika = 10/10 = 1 km/min
 
now speed of  rumpa = 20/12 = 1.667km/min
 
hence rumpa runs faster